the product of two consecutive integers is 55 more than their sum. Find the integers.
Answer by JimboP1977(311)

x(x+1)-55 = x+x+1
x^2+x-55 = 2x+1
x^2-x-56 = 0
(x-(1/2))^2 - (1/4) = 56
x = (225/4)^(1/2) + (1/2)
x = 8
So the two numbers are 8 and 9.